OpenStudy (anonymous):

The sum of the measures of 2 complimentary angles is 90 degrees. If one angle measures 27 degrees more than twice the measure of the other, find the measure of the smaller angle.

OpenStudy (anonymous):

The angles add to 90 _________ +__________ = 90 The first angle is unknown: x The second angle is 27 + 2 x Therefore we have x + 27 + 2x = 90 or 3x + 27 = 90 subtract 27 from both sides 3x = 63 or x = 21 The second angle is 27 + 2(21) = 69